I know I make common mistakes when solving equations can you help with the following:
The equation: 2(x -2) =x+3
First step in solving: 2x-2 =x+3
Can you tell me a clear explanation of what error has been made. What could be done to avoid this error?

Answer by AnlytcPhil(1806)

I know I make common mistakes when solving 
equations can you help with the following: 
The equation: 2(x -2) =x+3 
First step in solving: 2x-2 =x+3 
Can you tell me a clear explanation of what 
error has been made. What could be done to 
avoid this error?

Your error is in the distributive law or 
principle. You have

2(x - 2) = x + 3

Then you have

2x - 2 = x + 3

That is wrong. It should be

2x - 4 = x + 3

That's because when you have 2(x - 2) you multiply 
the red 2 first times the x, which you did,  but ALSO
you must multiply the red 2 times the blue - 2 as well.
What you did was to multiply the red 2 by the x and
then just copy down the blue - 2 without multiplying 
the red 2 by it to get - 4.

Edwin
AnlytcPhil@aol.com